OpenStudy (anonymous):

help me factorize 5y^2-9y-3 answer choices: A. 2.09, –0.29 B. 15, –13.2 C. 4.17, –0.57 D. 0.29, –2.09

OpenStudy (mathstudent55):

Are you familiar with the quadratic formula?

OpenStudy (mathstudent55):

\(ax^2 + bx + c = 0\) has solutions \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

You're welcome.